Required information [The following information applies to the questions displayed below.] Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March): Molding Fabrication Total Estimated total machine-hours used 2,500 1,500 4,000 Estimated total fixed manufacturing overhead $ 14,500 $ 17,700 $ 32,200 Estimated variable manufacturing overhead per machine-hour $ 3.20 $ 4.00 Job P Job Q Direct materials $ 31,000 $ 17,000 Direct labor cost $ 35,400 $ 14,700 Actual machine-hours used: Molding 3,500 2,600 Fabrication 2,400 2,700 Total 5,900 5,300 Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month. Required: For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1. What were the company’s predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your answers to 2 decimal places.)

Molding Department: ___ per MH

Fabrication Department: ___ per MH

2. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

Manufacturing Overhead Applied Job P:

Manufacturing Overhead Applied Job Q:

3. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)

Unit Product Cost:

4. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

Unit Product Cost:

5.

Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

Total Price for the Job P:

Total Price for the Job Q:

Selling Price per Unit Job P:

Selling Price per Unit job Q:

6. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

​​​​​​​Cost of goods sold:
Answer #1

1) Predetermine overhead rate = (2500*3.2+1500*4+32200)/4000 = 11.55 per MH

Molding department rate = 11.55 per MH

Fabrication department rate = 11.55 per MH

2) Manufacturing overhead applied to Job P = 5900*11.55 = 68145

Manufacturing overhead applied to Job Q = 5300*11.55 = 61215

3) Unit product cost of Job P = (31000+35400+68145)/20 = 6727

4) Unit product cost of Job Q = (17000+14700+61215)/30 = 3097

5) Calculate selling price and selling price per unit

Job P Job Q
Direct material 31000 17000
Direct labor 35400 14700
Overhead 68145 61215
Total manufacturing cost 134545 92915
Markup 107636 74332
Selling price 242181 167247
Selling price per unit 12109 5575

6) Cost of goods sold = 134545+92915 = $227460